# Quantum Fisher information analysis for absorption measurements with undetected photons

### v1

@@ I Introduction
Recent work has demonstrated that when combined with photon counting, SU(1,1) interferometers provide an unconditional quantum metrological advantage over conventional Mach–Zehnder interferometers [40]. From an experimental point of view, SU(1,1) interferometers have become the central platform for broadband quantum spectroscopy [24, 28, 29, 27, 17, 35].
Two additional experimental configurations have been developed to realize measurements with undetected photons. The induced-coherence (IC) scheme [44] closely resembles the SU(1,1) configuration but differs in that only the idler from the first source propagates through the second. A new signal photon is generated there that carries the encoded information from the idler. Interference between the two signal modes on a beam splitter then reveals the sample-dependent changes. The IC configuration enabled the first experimental realization of measurements with undetected photons [23] and has recently attracted renewed attention [8, 37].
@@ II.1 Lossless continuous wave model
$\frac{\partial}{\partial z}\hat{a}_{S}(\omega,z)=i\Delta k_{S}(\omega)\hat{a}_{S}(\omega,z)+i\gamma\hat{a}^{\dagger}_{I}(-\omega,z),$ (3)
$\frac{\partial}{\partial z}\hat{a}^{\dagger}_{I}(\omega,z)=-i\Delta k_{I}(\omega)\hat{a}^{\dagger}_{I}(\omega,z)-i\gamma^{*}\hat{a}_{S}(-\omega,z).$ (4)
where $\Delta k_{S/I}(\omega)=k_{S/I}(\omega)-\bar{k}_{S/I}$ are the dispersion relations detuned from their central wavevectors, $\gamma=|\gamma|e^{-i\Phi_{P}}$ represents the interaction strength which depends on the phase of the pump and many other material parameters [34, 33, 14], $x^{*}$ denotes the complex conjugate of $x$. In all the equations in this manuscript the variable $\omega$ is used to indicate a detuning from a central frequency. Thus, for example, $\hat{a}_{S}(\omega,z)$ denotes a Heisenberg evolved operator that destroys photons of frequency $\bar{\omega}_{S}+\omega$; similarly, $k_{I}(-\omega)$, describes the wavevector associated with frequency $\bar{\omega}_{I}-\omega$. This notation is necessary to avoid unnecessary cluttered equations. The equations of motion thus link the signal annihilation operator at frequency $\bar{\omega}_{S}+\omega$ to the idler creation operator at $\bar{\omega}_{I}-\omega$, a consequence of energy conservation, see Eq. (1).
@@ III SU(1,1) Model
For the SU(1,1) model, we treat the sources as ideal squeezers without loss and allow for the modes to incur loss and additional dispersion in between these ideal sources. We derive expressions for the moments in full generality and specify the case of idler only loss after. From the previous section, we know how the second order moments are modified when passing through the nonlinear region for both vacuum and general inputs. We now need to find expressions for how the moments are modified by additional dispersion and loss. To do so, we treat the loss and additional dispersion via beamsplitter interactions (with the bath modes taken to be Markovian quantum noise). In terms of the moments before the additional dispersion and loss (in) we find that the output moments (out) are given by
@@ VI.2.1 Induced coherence: access to ancilla and signal mode
where the superscript “TM” refers to the fact that we are considering two-modes and $H^{IC}_{\kappa_{I}}$ is the QFI with access to all modes (Eq. (54)). In Fig. 2(d) we show the two-mode QFI w.r.t. $\kappa_{I}$ for the IC model for different levels of gain. From this expression and by comparing the figures, we clearly see that the two-mode IC QFI is less than the full mode QFI. This is to be expected since we are removing information when tracing out the idler. An interesting question to investigate is to see whether or not the addition of the ancilla mode gives an estimation advantage when comparing to the single-mode SU(1,1) and DL models.
@@ VI.2.2 Single-mode operation
This tells us that the single-mode QFI is exactly the inverse of error propagation given by a signal intensity measurement for thermal Gaussian states. Therefore, if we only have access to a single mode, an intensity measurement will saturate the quantum Cramér-Rao bound. In Figs. 2(g-i) we show the single-mode QFI for the SU(1,1), DL, and IC models respectively. Note that for the single-mode IC configuration, where we choose phases such that $\sin(\frac{\Delta KL}{2}+\Phi)=1$ in Eq. (IV), we find that the $N^{\text{IC,BBS}}_{-}$ gives rise to the optimal QFI.

### v2

@@ I Introduction
Recent work has demonstrated that when combined with photon counting, SU(1,1) interferometers provide an unconditional quantum metrological advantage over conventional Mach–Zehnder interferometers [43]. In the single photon-pair generation regime, it has been shown that SU(1,1) interferometers combined with intensity difference measurements can provide a loss-tolerant advantage for absorption sensitivity compared to classical schemes [30]. From an experimental point of view, SU(1,1) interferometers have become the central platform for broadband quantum spectroscopy [26, 31, 32, 29, 19, 38].
Two additional experimental configurations have been developed to realize measurements with undetected photons. The induced-coherence (IC) scheme [49] closely resembles the SU(1,1) configuration but differs in that only the idler from the first source propagates through the second. A new signal photon is generated there that carries the encoded information from the idler. Interference between the two signal modes on a beam splitter then reveals the sample-dependent changes. The IC configuration enabled the first experimental realization of measurements with undetected photons [25] and has recently attracted renewed attention [8, 40, 47].
@@ II.1 Lossless continuous wave model
$\frac{\partial}{\partial z}\hat{c}_{S}(\omega,z)=i\Delta k_{S}(\omega)\hat{c}_{S}(\omega,z)+i\gamma\hat{c}^{\dagger}_{I}(-\omega,z),$ (3)
$\frac{\partial}{\partial z}\hat{c}^{\dagger}_{I}(\omega,z)=-i\Delta k_{I}(\omega)\hat{c}^{\dagger}_{I}(\omega,z)-i\gamma^{*}\hat{c}_{S}(-\omega,z).$ (4)
where $\Delta k_{S/I}(\omega)=k_{S/I}(\omega)-\bar{k}_{S/I}$ are the dispersion relations detuned from their central wavevectors, $\gamma=|\gamma|e^{-i\Phi_{P}}$ represents the interaction strength which depends on the phase of the pump and many other material parameters [37, 36, 15, 13], $x^{*}$ denotes the complex conjugate of $x$. In all the equations in this manuscript the variable $\omega$ is used to indicate a detuning from a central frequency. Thus, for example, $\hat{a}_{S}(\omega,z)$ denotes a Heisenberg evolved operator that destroys photons of frequency $\bar{\omega}_{S}+\omega$; similarly, $k_{I}(-\omega)$, describes the wavevector associated with frequency $\bar{\omega}_{I}-\omega$. This notation is necessary to avoid unnecessary cluttered equations. The equations of motion thus link the signal annihilation operator at frequency $\bar{\omega}_{S}+\omega$ to the idler creation operator at $\bar{\omega}_{I}-\omega$, a consequence of energy conservation, see Eq. (1).
For numerical evaluation, we discretize frequency space such that $\omega_{n}=\omega_{0}+n\Delta\omega|^{N-1}_{0}$ for an $N$-size grid and introduce new, dimensionless operators $\hat{a}_{S/I}(\omega_{n})=\hat{c}_{S/I}(\omega_{n})\sqrt{\Delta\omega}$. These new operators obey the same Heisenberg equations of motion (Eqs.(3) and (4))
@@ III SU(1,1) Model
For the SU(1,1) model, we treat the sources as ideal squeezers without loss and allow for the modes to incur loss and additional dispersion in between these ideal sources. Furthermore, we consider the case where the two squeezers are perfectly mode matched [18], i.e. where the input and output modes of the squeezers are identical. Misalignment of the squeezers could lead to distinguishability between the modes of the two squeezers which requires a completely different analysis [48] which is outside the scope of this manuscript. We derive expressions for the moments in full generality and specify the case of idler only loss after. From the previous section, we know how the second order moments are modified when passing through the nonlinear region for both vacuum and general inputs. We now need to find expressions for how the moments are modified by additional dispersion and loss. To do so, we treat the loss and additional dispersion via beamsplitter interactions (with the bath modes taken to be Markovian quantum noise). In terms of the moments before the additional dispersion and loss (in) we find that the output moments (out) are given by
@@ VI.2.1 Induced coherence: access to ancilla and signal mode
where the superscript “TM” refers to the fact that we are considering two-modes and $H^{IC}_{\kappa_{I}}$ is the QFI with access to all modes (Eq. (54)). This result agrees with those of [47]. In Fig. 2(d) we show the two-mode QFI w.r.t. $\kappa_{I}$ for the IC model for different levels of gain. From this expression and by comparing the figures, we clearly see that the two-mode IC QFI is less than the full mode QFI. This is to be expected since we are removing information when tracing out the idler. An interesting question to investigate is to see whether or not the addition of the ancilla mode gives an estimation advantage when comparing to the single-mode SU(1,1) and DL models.
@@ VI.2.2 Single-mode operation
This tells us that the single-mode QFI is exactly the inverse of error propagation given by a signal intensity measurement for thermal Gaussian states. Therefore, if we only have access to a single mode, an intensity measurement will saturate the quantum Cramér-Rao bound. In Figs. 2(g-i) we show the single-mode QFI for the SU(1,1), DL, and IC models respectively. Note that for the single-mode IC configuration, where we choose phases such that $\sin(\frac{\Delta KL}{2}-\Phi)=1$ in Eq. (IV), we find that the $N^{\text{IC,BBS}}_{-}$ gives rise to the optimal QFI.